# Tuning-free Visual Effect Transfer across Videos

### v3

@@ 1 Introduction
Generative models have enabled the automatic synthesis and manipulation of images and videos with remarkable realism and diversity [blattmann2023stable, wan2025wan, kong2024hunyuanvideo, hacohen2024ltx, hong2022cogvideo, polyak2024movie, rombach2022high, podell2023sdxl, esser2024scaling, flux2024]. Recent advances in video generation have expanded user control, allowing video editing via text prompts, keyframes, or depth maps [wan2025wan, jiang2025vace, zhao2025controlvideo]. However, most models focus on semantic edits such as modifying objects [zi2025se, polyak2024movie], scenes [zi2025se, polyak2024movie], or styles [liu2023stylecrafter, ye2025stylemaster, zi2025se], while overlooking “temporal effects” (see Figure  1): effects that unfold over time and define the emotional, stylistic, and cinematic character of a video. Effects such as dynamic lighting, intricate camera movements, or character transformations remain difficult to express through text or current visual inputs.
@@
While reference-based editing has shown promise for images [ye2023ip, kumari2023multi, ruiz2023dreambooth], extending it to videos poses unique challenges. First, building a dataset is non-trivial: image-to-video (I2V) effects require consistent triplets in which a single static input produces multiple effect-rich outputs, and video-to-video (V2V) effects demand motion-consistent transformations that change only the temporal effect. Without such data, existing systems may fail to learn how to transfer temporal effects independently of the reference video’s content or motion. Second, the model must extract temporal dynamics from the reference and seamlessly integrate them with the target’s motion and appearance to produce coherent, high quality results. This requires disentangling the effect itself from the specific content and motion of the reference.
We introduce a reference-based video generation dataset and a training method dubbed RefVFX that transfers complex temporal effects from a reference video onto a target image or video in a feed-forward manner. Built upon recent text-to-video diffusion backbones [wan2025wan, hacohen2024ltx, kong2024hunyuanvideo], our model jointly conditions on three inputs: a reference video that provides the temporal effect, an input image or video that defines the scene content and motion, and a text prompt that offers high-level semantic guidance. Through this conditioning, RefVFX learns to integrate the reference’s temporal dynamics harmoniously with the input’s appearance and motion, producing coherent and visually consistent results.
@@
We evaluate RefVFX across diverse unseen temporal effects, comparing it with recent prompt-based and reference-guided video editing models [wan2025wan, jiang2025vace, Hleihil2025DecartAI]. Qualitatively, our method produces coherent, visually consistent videos that successfully integrate the transferred effects while preserving the motion and appearance of the input. Quantitatively, RefVFX achieves higher scores in reference video embedding similarity to the reference effect video; however, existing metrics only partially reflect the aesthetic and temporal nuances of effect transfer. To better assess perceptual quality, we conduct a human preference study, where participants consistently favor our results, highlighting the importance of human judgment for evaluating dynamic visual effects. Notably, the model generalizes to unseen categories of effects and operates in a feed-forward manner without optimization at inference time, demonstrating both robustness and efficiency. Our code will be released upon publication.
@@ 2 Related Work
Text-to-Video Generation. Recent advances in text-to-video generation have significantly improved video quality, temporal coherence, and prompt following [blattmann2023stable, wan2025wan, kong2024hunyuanvideo, hacohen2024ltx, hong2022cogvideo, polyak2024movie, ma2024latte, cloneofsimo2024auraflow, ho2022imagen, ho2022video]. Similar to text-to-image generation [rombach2022high, podell2023sdxl, esser2024scaling, flux2024], these models are largely based on diffusion or flow-matching in latent space [song2020denoising, dhariwal2021diffusion, lipman2022flow, albergo2023stochastic, liu2022flow]. Early work adopted U-Net backbones [ronneberger2015u, singer2022make, guo2023animatediff], while modern approaches employ diffusion transformers [peebles2023scalable] that embed videos into latent patches and apply bidirectional attention with text [wan2025wan, hacohen2024ltx, polyak2024movie, ho2022imagen, hong2022cogvideo, kong2024hunyuanvideo]. Beyond pure text conditioning, some models perform image-to-video or first–last-frame generation [wan2025wan, hacohen2024ltx], where a user provides keyframes and a text prompt to guide synthesis. However, these settings modify only edge frames, whereas our approach conditions on both an entire input video and a reference effect video, enabling transfer of temporal effects that evolve throughout the sequence while preserving input motion and appearance.
@@
Reference-Based Controllable Generation. Reference-based conditioning uses a reference image or video that is not spatially aligned with the output. In text-to-image models, it is widely used for identity preservation [kumari2025generating, chen2024anydoor, li2023blip, parmar2025object, song2024moma, wei2023elite, xiao2025fastcomposer, guo2024pulid, ye2023ip, patashnik2025nested, ruiz2023dreambooth] or style transfer [huang2023composer, hertz2024style, rout2024rb, gao2025styleshot, xing2024csgo, zhang2024finestyle, sohn2023styledrop]. These signals are injected via per-reference optimization [sohn2023styledrop, gao2025styleshot, ruiz2023dreambooth], cross-attention [li2023blip, parmar2025object, guo2024pulid, ye2023ip, song2024moma, xing2024csgo], or added reference tokens [kumari2025generating, hertz2024style, rout2024rb]. Video generative models extend these paradigms temporally [wang2023videocomposer, liu2025phantom, liu2023stylecrafter, he2024id, yuan2025identity, huang2025conceptmaster, liang2025movie, chen2024videoalchemy, ye2025stylemaster], typically conditioning on one or more static reference images (identity or style) alongside text. Such methods maintain consistent appearance but cannot model temporally evolving effects. More recent finetuning based methods such as Dynamic Concepts [dynamic_concepts, abdal2025zeroshotdynamicconceptpersonalization] are also inefficient, requiring either a new LoRA per effect instance or trained on limited data hampering scalability. In contrast, RefVFX conditions on a reference video that encodes dynamic, time-varying effects, allowing generalizable transfer of phenomena such as lighting changes, camera motion, stylistic transitions or identity-based conditioning.
@@
Video Editing. Video editing has progressed through both zero-shot and supervised methods [brooks2023instructpix2pix, labs2025flux, wu2025qwen, wei2024omniedit, hertz2022prompt, tumanyan2023plug, cao2023masactrl, zhang2023magicbrush, kulikov2025flowedit, brack2024ledits++, deutch2024turboedit]. Pretrained text-to-image models can perform zero-shot edits [hertz2022prompt, tumanyan2023plug, cao2023masactrl, kulikov2025flowedit, brack2024ledits++, deutch2024turboedit], while paired datasets of object or style edits [wei2024omniedit, boesel2024improving] have enabled general-purpose editing models [brooks2023instructpix2pix, labs2025flux, wu2025qwen]. Video editing methods follow similar patterns: some are zero-shot [ceylan2023pix2video, geyer2023tokenflow], while others train paired text-driven editors [cheng2023consistent, zi2025se, liu2025generative]. However, existing methods rely solely on text and an input video, enabling semantic or appearance edits but offering no mechanism to control how visual properties evolve over time. Our approach introduces an additional reference video encoding a temporally varying effect, enabling control over dynamic phenomena such as lighting transitions, motion-dependent effects, and stylistic evolution that static, per-frame, or text-only paradigms cannot capture.
@@
Figure 2: Dataset Examples. We show example triplets from each of our datasets. (a) We curate a reference video + image to video dataset by collecting Low Rank Adapters (LoRAs) [hu2022lora] for different Image to Video effects online. For each effect, we can apply its corresponding LoRA to two separate images to create a triplet. (b) We create a custom pipeline for generating text guided reference video + input video to video effects. For more details, see Figure 4. (c) We generate a large scale set of (ref video, input video, output video) triplets by curating specific code pipelines that apply specific, detailed effects to arbitrary videos. Armed with a specific code based effect and a fixed set of hyperparameters, we can apply the exact effect to an arbitrary number of input videos to create many triplets.
@@ 3.1 Overview
Our goal is to transfer a temporally evolving visual effect from a reference video onto a separate input image or video. The resulting output should preserve the content and motion of the input while exhibiting the temporal dynamics shown in the reference. To achieve this, we introduce RefVFX, which combines (1) a large-scale effect–aligned triplet dataset and (2) a diffusion-based video generation model conditioned jointly on the reference video, input video, and text.
@@ 3.2 Training Data
3.2.1 Image to Video Data:
We curate a set of image-to-video effects using open-source Low-Rank Adapters (LoRAs) [hu2022lora] trained on small sets of videos sharing a common visual effect atop a base image-to-video model [wan2025wan]. Each adapter is treated as a distinct effect $E_{i}$. For every effect, we generate a large collection of videos from diverse input images synthesized via high-quality image generation models [wu2025qwen, labs2025flux], where each per-adapter set forms one effect category. Triplets are composed of a reference video with effect $E_{i}$, the first frame of the target, and the corresponding output video. In total, we source 43 LoRAs with an average of 300 videos per effect, yielding over 14K video clips. See Fig. 3 for example LoRAs.
3.2.2 Video to Video Data:
@@
Given an arbitrary effect prompt, our method produces paired videos $(V,V^{\prime})$, consisting of an original video $V$ and its edited counterpart $V^{\prime}$ exhibiting effect $E_{i}$ (Fig. 4, 3). Scalable V2V Algorithm. The process proceeds for generating these $(V,V^{\prime})$ pairs is as follows: We begin by generating a subject image $I$ using a high-quality image generation model [flux2024, wu2025qwen]. Using an image editing model [labs2025flux, wu2025qwen], we modify $I$ to produce an image $I^{\prime}$ that depicts the same subject under a new pose, camera angle, and facial expression, which will be used to anchor the last frame of our video pair. To obtain the last frame of the edited video with effect $E_{i}$, we apply an image editing model on $I^{\prime}$ [labs2025flux, wu2025qwen] to apply effect $E_{i}$, resulting in $I^{\prime\prime}$. Given the unedited image $I$ and the edited image $I^{\prime}$, we synthesize video $V$ using a first–last frame interpolation model [wan2025wan]. Once we have $V$, we extract intermediate poses using an image-to-pose model [yang2023effective], which will be used as conditioning for our second video generation. Finally, we generate the edited video $V^{\prime}$ by leveraging a conditional video generation model [jiang2025vace], which can take in independent conditioning signals for each output video frame generated. We take advantage of this property, and condition the video generation on:
@@
This enables framewise consistency between the input and output video as well as realistic temporal transitions between the unedited initial frame $I$ and stylized final frame $I^{\prime\prime}$. For the a visual, see Figure 4.
@@
To ensure diversity, we automatically generate a large set of motion-based effect prompts using GPT-4o [hurst2024gpt]. These effects span multiple semantic and visual categories, which we detail in the supplement.
3.2.3 Synthetic Dataset Creation.
We further construct a large-scale dataset of synthetic motion-based effects through programmatic generation. Our approach enables scalable and reproducible creation of diverse video effects with fine-grained control over appearance and temporal dynamics.
Base Videos. We source input videos from the Senorita dataset [zi2025se], which provides high-quality clips with accompanying foreground–background segmentation masks obtained using SegmentAnything [ravi2024sam]. These masks allow effects to be applied selectively to the full frame, the foreground, or the background.
@@ 3.3 Model Architecture and Conditioning
We build upon the Wan conditioning model architecture [wan2025wan, jiang2025vace], which extends text-to-video diffusion models to image-to-video (I2V), first-last frame to video (FLF2V), depth to video, and other conditioning types by concatenating noisy latent channels with conditional and mask channels. We extend this mechanism to jointly condition on both a reference video effect and an input video.
@@ 3.4 Implementation Details
We fine-tune the Wan [wan2025wan] 14 Billion parameter First-Last Frame to Video model with Low Rank Adapters for 10K steps with a batch size of 8 on a single node of 8 NVIDIA A100 GPUs. We sample from each of the three datasets equally during training, and drop the reference effect video conditioning, control video conditioning, and text prompts with low probability to allow for various forms of classifier-free guidance [ho2022classifier, brooks2023instructpix2pix] during inference time. We also add the true last frame from the target video to the model with low probability to allow for the first and last frame capabilities of the underlying model to be retained under our new setup.
@@
Baseline RVA IVA RVA IVA
Lucy Edit 94.3 / +1.26 98.2 / +1.49
VACE (Depth) 91.7 / +1.07 70.0 / +0.05 98.3 / +1.51 Tie / +0.00
VACE (Pose) 78.3 / +0.72 90.0 / +0.11 96.7 / +1.48 Tie / +0.00
No Ref 92.9 / +0.18 Tie / +0.00 93.6 / +0.94 Tie / +0.00
Table 1: User Study Results for Neural V2V and Code-Based Edits. Each cell shows Win Rate (%) / Mean Score ($\pm$2 scale). Our method shows large preference over base models in Reference Video Adherence (RVA), and gains in Input Video Adherence (IVA)
Baseline Reference Video Adherence
Wan2.1 74.7 / +0.57
VACE (I2V) 80.0 / +0.69
No Ref 81.5 / +0.15
Table 2: User Study Results for I2V. Each cell shows Win Rate (%) / Mean Score ($\pm$2 scale). Our method shows large user preference over baselines in Reference Video Adherence.
@@ 4.1 Validation Datasets
Video-to-Video Testing. We generate a validation set of video-to-video pairs using the pipeline introduced in Section 3.2.2, with novel prompts not seen during training. Each resulting pair (reference video, input video) is manually filtered to ensure high perceptual quality.To further test cross-domain generalization, we reuse the unseen LoRA effects from the previous section but replace the input images with novel input videos, producing over 100 test pairs in total. Finally, we synthesize an additional set of temporally varying effects using the programmatic procedure from Section 3.2.3. Although these effects share the same families of operations and transitions as those used for training, their sampled hyperparameters are unseen at train time.

### v4

@@ 1 Introduction
Generative models have enabled the automatic synthesis and manipulation of images and videos with remarkable realism and diversity [5, 74, 38, 22, 30, 58, 62, 57, 17, 43]. Recent advances in video generation have expanded user control, allowing video editing via text prompts, keyframes, or depth maps [74, 36, 88]. However, most models focus on semantic edits such as modifying objects [89, 58], scenes [89, 58], or styles [49, 83, 89], while overlooking “temporal effects” (see Figure  1): effects that unfold over time and define the emotional, stylistic, and cinematic character of a video. Effects such as dynamic lighting, intricate camera movements, or character transformations remain difficult to express through text or current visual inputs.
@@
While reference-based editing has shown promise for images [82, 40, 65], extending it to videos poses unique challenges. First, building a dataset is non-trivial: image-to-video (I2V) effects require consistent triplets in which a single static input produces multiple effect-rich outputs, and video-to-video (V2V) effects demand motion-consistent transformations that change only the temporal effect. Without such data, existing systems may fail to learn how to transfer temporal effects independently of the reference video’s content or motion. Second, the model must extract temporal dynamics from the reference and seamlessly integrate them with the target’s motion and appearance to produce coherent, high quality results. This requires disentangling the effect itself from the specific content and motion of the reference.
We introduce a reference-based video generation dataset, training method, and final framework of finetuned model along with an updated conditioning mechanism dubbed RefVFX that transfers complex temporal effects from a reference video onto a target image or video in a feed-forward manner. Built upon recent text-to-video diffusion backbones [74, 22, 38], our model jointly conditions on three inputs: a reference video that provides the temporal effect, an input image or video that defines the scene content and motion, and a text prompt that offers high-level semantic guidance. Through this conditioning, RefVFX learns to integrate the reference’s temporal dynamics harmoniously with the input’s appearance and motion, producing coherent and visually consistent results.
@@
We evaluate RefVFX across diverse unseen temporal effects, comparing it with recent prompt-based and reference-guided video editing models [74, 36, 26]. Qualitatively, our method produces coherent, visually consistent videos that successfully integrate the transferred effects while preserving the motion and appearance of the input. Quantitatively, RefVFX achieves higher scores in reference video embedding similarity to the reference effect video; however, existing metrics only partially reflect the aesthetic and temporal nuances of effect transfer. To better assess perceptual quality, we conduct a human preference study, where participants consistently favor our results, highlighting the importance of human judgment for evaluating dynamic visual effects. Notably, the model generalizes to unseen categories of effects and operates in a feed-forward manner without optimization at inference time, demonstrating both robustness and efficiency. Our code will be released upon publication.
@@ 2 Related Work
Text-to-Video Generation. Recent advances in text-to-video generation have significantly improved video quality, temporal coherence, and prompt following [5, 74, 38, 22, 30, 58, 53, 14, 28, 29]. Similar to text-to-image generation [62, 57, 17, 43], these models are largely based on diffusion or flow-matching in latent space [69, 16, 48, 4, 52]. Early work adopted U-Net backbones [63, 67, 20], while modern approaches employ diffusion transformers [56] that embed videos into latent patches and apply bidirectional attention with text [74, 22, 58, 28, 30, 38]. Beyond pure text conditioning, some models perform image-to-video or first–last-frame generation [74, 22], where a user provides keyframes and a text prompt to guide synthesis. However, these settings modify only edge frames, whereas our approach conditions on both an entire input video and a reference effect video, enabling transfer of temporal effects that evolve throughout the sequence while preserving input motion and appearance.
@@
Reference-Based Controllable Generation. Reference-based conditioning uses a reference image or video that is not spatially aligned with the output. In text-to-image models, it is widely used for identity preservation [42, 12, 46, 54, 70, 77, 79, 21, 82, 55, 65] or style transfer [32, 25, 64, 18, 80, 85, 68]. These signals are injected via per-reference optimization [68, 18, 65], cross-attention [46, 54, 21, 82, 70, 80], or added reference tokens [42, 25, 64]. Video generative models extend these paradigms temporally [75, 50, 49, 23, 84, 33, 47, 11, 83], typically conditioning on one or more static reference images (identity or style) alongside text. Such methods maintain consistent appearance but cannot model temporally evolving effects. More recent finetuning based methods such as Dynamic Concepts [3, 2] are also inefficient, requiring either a new LoRA per effect instance or trained on limited data hampering scalability. In contrast, RefVFX conditions on a reference video that encodes dynamic, time-varying effects, allowing generalizable transfer of phenomena such as lighting changes, camera motion, stylistic transitions or identity-based conditioning.
@@
Video Editing. Video editing has progressed through both zero-shot and supervised methods [8, 44, 78, 76, 24, 73, 9, 86, 39, 7, 15]. Pretrained text-to-image models can perform zero-shot edits [24, 73, 9, 39, 7, 15], while paired datasets of object or style edits [76, 6] have enabled general-purpose editing models [8, 44, 78]. Video editing methods follow similar patterns: some are zero-shot [10, 19], while others train paired text-driven editors [13, 89, 51]. However, existing methods rely solely on text and an input video, enabling semantic or appearance edits but offering no mechanism to control how visual properties evolve over time. Our approach introduces an additional reference video encoding a temporally varying effect, enabling control over dynamic phenomena such as lighting transitions, motion-dependent effects, and stylistic evolution that static, per-frame, or text-only paradigms cannot capture.
@@
Figure 2: Dataset Examples. We show example triplets from each of our datasets. (a) We curate a reference video + image to video dataset by collecting Low Rank Adapters (LoRAs) [31] for different Image to Video effects online. For each effect, we can apply its corresponding LoRA to two separate images to create a triplet. (b) We create a custom pipeline for generating text guided reference video + input video to video effects. For more details, see Figure 4. (c) We generate a large scale set of (ref video, input video, output video) triplets by curating specific code pipelines that apply specific, detailed effects to arbitrary videos. Armed with a specific code based effect and a fixed set of hyperparameters, we can apply the exact effect to an arbitrary number of input videos to create many triplets.
@@ 3.1 Overview
Our goal is to transfer a temporally evolving visual effect from a reference video onto a separate input image or video. The resulting output should preserve the content and motion of the input while exhibiting the temporal dynamics shown in the reference. To achieve this, we introduce RefVFX, which combines (1) a large-scale effect–aligned triplet training dataset, (2) a pretrained diffusion-based video generation model, and (3) an updated conditioning scheme allowing the video generation model to condition jointly on the reference video, input video, and text.
@@ 3.2 Training Data
3.2.1 Image to Video Effects:
We curate a set of image-to-video effects using open-source Low-Rank Adapters (LoRAs) [31] trained on small sets of videos sharing a common visual effect atop a base image-to-video model [74]. Each adapter is treated as a distinct effect $E_{i}$. For every effect, we generate a large collection of videos from diverse input images synthesized via high-quality image generation models [78, 44], where each per-adapter set forms one effect category. Triplets are composed of a reference video with effect $E_{i}$, the first frame of the target, and the corresponding output video. In total, we source 43 LoRAs with an average of 300 videos per effect, yielding over 14K video clips. See Fig. 3 for example LoRAs.
3.2.2 Video to Video Effects:
@@
Given an arbitrary effect prompt, our method produces paired videos $(V,V^{\prime})$, consisting of an original video $V$ and its edited counterpart $V^{\prime}$ exhibiting effect $E_{i}$ (Fig. 4, 3). Scalable V2V Algorithm. The process for generating these $(V,V^{\prime})$ pairs is as follows: We begin by generating a subject image $I$ using a high-quality image generation model [43, 78]. Using an image editing model [44, 78], we modify $I$ to produce an image $I^{\prime}$ that depicts the same subject under a new pose, camera angle, and facial expression, which will be used to anchor the last frame of our video pair. To obtain the last frame of the edited video with effect $E_{i}$, we apply an image editing model on $I^{\prime}$ [44, 78] to apply effect $E_{i}$, resulting in $I^{\prime\prime}$. Given the unedited image $I$ and the edited image $I^{\prime}$, we synthesize video $V$ using a first–last frame interpolation model [74]. Once we have $V$, we extract intermediate poses using an image-to-pose model [81], which will be used as conditioning for our second video generation. Finally, we generate the edited video $V^{\prime}$ by leveraging a conditional video generation model [36], which can take in independent conditioning signals for each output video frame generated. We take advantage of this property, and condition the video generation on:
@@
This enables framewise consistency between the input and output video as well as realistic temporal transitions between the unedited initial frame $I$ and stylized final frame $I^{\prime\prime}$. For a visual, see Figure 4.
@@
To ensure diversity, we automatically generate a large set of motion-based effect prompts using GPT-4o [35]. These effects span multiple semantic and visual categories, which we detail in the supplement.
3.2.3 Programmatic Temporal Effects.
We further construct a large-scale dataset of programmatically defined temporal effects applied to real videos. Our approach enables scalable and reproducible creation of diverse video effects with fine-grained control over appearance and temporal dynamics.
Base Videos. We source input videos from the Senorita dataset [89], which provides high-quality clips with accompanying foreground–background segmentation masks obtained using SegmentAnything [61]. These masks allow effects to be applied selectively to the full frame, the foreground, or the background.
@@ 3.3 Model Architecture and Conditioning
We build upon the Wan2.1 conditioning model architecture [74, 36], which extends text-to-video diffusion models to image-to-video (I2V), first-last frame to video (FLF2V), depth to video, and other conditioning types by concatenating noisy latent channels with conditional and mask channels. We extend this mechanism to jointly condition on both a reference video effect and an input video.
@@ 3.4 Implementation Details
We fine-tune the Wan2.1 [74] 14 Billion parameter First-Last Frame to Video model with Low Rank Adapters for 10K steps with a batch size of 8 on a single node of 8 NVIDIA A100 GPUs. We sample from each of the three datasets equally during training, and drop the reference effect video conditioning, control video conditioning, and text prompts with low probability to allow for various forms of classifier-free guidance [27, 8] during inference time. We also add the true last frame from the target video to the model with low probability to allow for the first and last frame capabilities of the underlying model to be retained under our new setup.
@@
Baseline RVA IVA OM RVA IVA OM
Lucy Edit 62.3 $\pm$ 2.9 60.7 $\pm$ 2.8 65.7 $\pm$ 3.1 64.7 $\pm$ 4.5 61.4 $\pm$ 5.0 65.3 $\pm$ 5.0
VACE (Depth) 58.3 $\pm$ 2.5 59.1 $\pm$ 2.6 66.8 $\pm$ 2.1 65.7 $\pm$ 5.4 56.2 $\pm$ 5.0 61.9 $\pm$ 5.0
VACE (Pose) 57.0 $\pm$ 2.1 53.7 $\pm$ 2.2 60.0 $\pm$ 2.2 62.9 $\pm$ 4.8 63.3 $\pm$ 4.9 64.3 $\pm$ 4.7
No Ref 57.5 $\pm$ 2.8 51.3 $\pm$ 3.1 67.2 $\pm$ 2.9 59.5 $\pm$ 3.5 54.3 $\pm$ 4.0 61.3 $\pm$ 3.9
Table 1: User Study Results for Neural V2V and Code-Based Edits. Each cell shows Win Rate (%) with standard deviation. Our method shows large preference over base models in Reference Video Adherence (RVA), Input Video Adherence (IVA), and Overall Match (OM), where we ask users to judge which method best adheres to the input video while applying the reference video effect.
Baseline Reference Video Adherence Overall Match
Wan2.1 57.1 $\pm$ 2.4 63.4 $\pm$ 2.4
VACE (I2V) 57.4 $\pm$ 3.2 62.9 $\pm$ 2.6
No Ref 56.8 $\pm$ 2.3 61.8 $\pm$ 3.4
Table 2: User Study Results for I2V. Each cell shows Win Rate (%) with standard deviation. Our method shows large user preference over baselines in Reference Video Adherence and Overall Match (OM), where we ask users to judge which method best preserves the input image while applying the reference video effect.
@@ 4.1 Validation Datasets
Video-to-Video Testing. We generate a validation set of Reference Video Effect + Input Video pairs for our reference effect based video to video task. For the first source of data, we reuse the unseen LoRA effects from the previous section as reference effect videos and replace the input images with unseen real input videos to produce the input video pair. Secondly, we use our pipeline introduced in Section 3.2.2 with novel prompts not seen during training for a portion of the validation dataset. Each resulting pair (reference video, input video) generated using our method is manually filtered to ensure high perceptual quality. Finally, we synthesize an additional set of temporally varying effects using the programmatic procedure from Section 3.2.3, but with hyperparameter choices unseen during training. We collect over 100 validation instances, and separate the Code-Based Video-to-Video Effects from the Neural Video-to-Video effects when reporting results.